OpenStudy (anonymous):

i have a question with the answer that i have to explain why that is the correct answer please explain to me why that answer is right .5. What is the probability that a point randomly chosen in the large triangle is also in Region A? answer is 75%

OpenStudy (anonymous):

here is the graph

OpenStudy (anonymous):

this is Geometric Probability

jimthompson5910 (jim_thompson5910):

what is the area of the larger triangle?

OpenStudy (anonymous):

it just shows the picture of the graph and some numbers on there it just saids the answer is 75% and to explain why is that

jimthompson5910 (jim_thompson5910):

are you able to find the area of the larger triangle?

OpenStudy (anonymous):

no can you help me please

jimthompson5910 (jim_thompson5910):

we have this triangle, agreed? 

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the area of this triangle is A = (b*h)/2 = (14*10)/2 = 140/2 = 70

jimthompson5910 (jim_thompson5910):

The smaller triangle looks like this 

jimthompson5910 (jim_thompson5910):

what is the area of this triangle?

OpenStudy (anonymous):

17.5?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so we divide the smaller triangle area by the larger triangle area 17.5/70 = 0.25 therefore, 25% of the area is just in the smaller triangle (and not in region A)

jimthompson5910 (jim_thompson5910):

so 100% - 25% = 75% must be in the remaining bit of area, aka region A

OpenStudy (anonymous):

oh okay i think i get it so first we had to find the areas for both triangles then divide the answers and got 25% take that away from a 100 and thats how we get 75%